OpenStudy (anonymous):

Eric can clean the car in 24 minutes. How long does it take Jackie to wash the car alone if Eric and Jackie can clean the car together in 6 minutes?

OpenStudy (raden):

Erick + Jackie = x : + :)

OpenStudy (anonymous):

so would i do 6x=24?

OpenStudy (anonymous):

im not really understanding what you mean?

OpenStudy (raden):

the formula if both work together is multiply the time of both divided by the sum of time of both, in this case is given

OpenStudy (raden):

let if Jackie work alone his time be x, so we have an equation : 24x/(24+x) = 6 solve for x

OpenStudy (anonymous):

I got x=18 is that right?

OpenStudy (raden):

how, show me ur steps :)

OpenStudy (anonymous):

nevermind i got -18 24+x=6 -24 -24 x=-18

OpenStudy (raden):

minus --"

OpenStudy (anonymous):

but it cant be -18 because it wouldnt make sense

OpenStudy (anonymous):

negative 18

OpenStudy (raden):